The length of a rectangle is 3 ft more than twice the width, and the area of the rectangle is 77 sq.ft. Find the dimensions of the rectangle
Answer by TimothyLamb(4379)

L = 2w + 3
Lw = 77
---
Lw = 77
(2w + 3)w = 77
2ww + 3w - 77 = 0
---
the above quadratic equation is in standard form, with a=2, b=3 and c=-77
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 3 -77
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 5.5
w = -7
---
the negative root doesn't fit the problem statement, so use the positive root:
w = 5.5
L = 2w + 3
L = 2*(5.5) + 3
L = 14
---
answer:
L = 14 ft
w = 5.5 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php